# Gibbs randomness-compression proposition: An efficient deep learning

### v2

@@ VI Numerical results
We also reported Gibbs entropy over train-compress iterative pruning in Figures 4 and 5. As complexity reduces, Gibbs entropy over measurement vectors also decreases. Pearson correlations between these and performance reduction are computed for DTC and random pruning. They read 0.9174 and 0.9412 respectively, empirically supporting our proposition.
Further study on the effect of frequency of dual tomographic procedure during train-compress iterations, variation of the resulting performances as we fixed the initialization for exact reproducibility, as estimating uncertainty reliably requires very large repeats, more optimizations on the penalty term of inverse compressed sensing procedure and the effect of binning on the Gibbs entropy on the measurement vectors appears to be good avenues to explore.
VII Conclusion
Contributions of the work appear to be quite stimulating for both theory and practice in three folds 1) we have formed mathematically more explicit relationship between directed randomness and complexity reduction compared to algorithmic randomness framework of SKC and demonstrated this empirically on our train-compress framework on the canonical deep learning task of classifying handwritten images. 2) formulation of an inverse compressed sensing framework, whereby measurements are generated rather than being a primary process 3) a new train-compress iterative pruning approach at neuronal level for deep learning is shown as a promising new approach for immensely energy efficient deep learning, supporting the lottery tickets hypothesis as well.
@@ Acknowledgment
Two Jupyter notebooks are provided for prototype implementations and reproducing the results presented in this work, named $dtcp\_analysis.ipynb$ and $dtcp\_generate\_data.ipynb$

### v3

@@ 6.1 Learning performance over compression cycles
6.2 Evidence for the proposition
We also reported Gibbs entropy over train-compress iterative pruning in Figures 4 and 5. As complexity reduces, Gibbs entropy over measurement vectors also decreases. Pearson correlations between these and performance reduction are computed for DTC and random pruning. They read 0.9174 and 0.9412 respectively, empirically supporting our proposition.This result demonstrated a strong empirical support for the Theorem 1, Gibbs-randomness compression proposition. While both DTC and random pruning have obviously random mechanism the Gibbs-compression proposition requires. As random pruning correlations is much higher, this strengthen the proposition’s statement. Moreover comonotonic behavior is also verified by the visual inspection between Figure 1 and 3 versus Figure 4 and 5.
6.3 Limitations and open improvements
We have developed foundations in developing unified view of distinct ideas of utilization of compressed sensing, dual-tomographic compression and the proposition that randomness and compression are highly correlated. Our experiments shown first promising steps supporting evidence for the proposition and utility of the DTC approach. We understand that more in-depth experimentation are required in advancing and supporting further these proposed methods and the proposition given in Theorem 1.
Further study on the effect of frequency of dual tomographic procedure during train-compress iterations, more optimizations on the penalty term of inverse compressed sensing procedure, the effect of binning on the Gibbs entropy on the measurement vectors and demonstrations with different architectures appears to be good avenues to explore for expanded future studies that requires much larger compute resources for advancing the approach with more diverse datasets.
7 Conclusion
There are multiple original contributions of the work appear to be quite stimulating for both theory and practice. 1) The formulation of an inverse compressed sensing framework, where measurements are generated rather than being a primary optimization process, 2) a new train-compress iterative pruning approach at neuronal level (layers reduced in one go, rather than pruning connections) for deep learning, via new technique called Dual-Tomographic Compression (DTC), is shown as a promising new approach for immensely energy efficient deep learning, 3) demonstration of accelerated lottery tickets hypothesis, that we can find sub-networks much faster during train-compress cycle rather that fully train and find the subnetwork, 4) we have formed mathematically more explicit relationship between directed randomness and complexity reduction compared to algorithmic randomness framework of SKC via a simple theorem, and demonstrated this empirically on our train-compress framework on the canonical deep learning task of classifying handwritten images as a promising results.
@@ Acknowledgment
Author is also grateful for constructive comments from the Neural Networks community, strengthen the paper’s technical case and presentation.
@@
Two Jupyter notebooks are provided for prototype implementations and reproducing the results presented in this work, named dtcp_analysis.ipynbdtcp\_analysis.ipynb and dtcp_generate_data.ipynbdtcp\_generate\_data.ipynb